OpenStudy (anonymous):

a solar thermal hot water system can off set 200 kW.h/month, so if your incoming temp is12c and your main usage temp is 42c how much hot water does that 200kW.h produce?

OpenStudy (anonymous):

Now we are talking. Let's note that \[\rm {kW \cdot hr \over month} \times {1 month \over 30 days} \times {1 day \over 24 hr} = {1 \over 720} kW\] Follow so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now, we can write our heat rate equation as follows. \[\dot Q \left [ \rm kW \right] = \dot m \left [kg \over s \right] \cdot c_p \left [ \rm J \over kg \cdot K \right] \cdot \Delta T \left [ \rm K \right]\]Note that we are left with the same units on the left and right hand sides because \(kW = {J \over s} \)

OpenStudy (anonymous):

We right this as\[{200 \over 720} = \dot m \cdot 4183 \cdot [42 - 12 ]\] If you wish to find the amount of water we can heat in a month, \[\dot m \left [ \rm kg \over month \right] = \dot m \left [ \rm kg \over s \right] \cdot \left [ \rm 3600 \sec \over 1 hour \right] \cdot \left [ \rm 720 hour \over month \right]\]

OpenStudy (anonymous):

*write